OpenStudy (aminamoazzam):

chuck cycles skyline drive . he cycles 60 km at an average speed of x km/h he then cycles a further 45 km at an average speed of (x+4) km/h. his total journey time is 6 hours . write doun an eqn in x and shoe that it implifies to 2x^2-27x-80=0.

OpenStudy (mathstudent55):

\(speed = \dfrac{distance}{time} \) \(time = \dfrac{distance}{speed} \) \(\dfrac{60}{x} + \dfrac{45}{x + 4} = 6\) \(\dfrac{60}{x}\times \dfrac{x+4}{x+4} + \dfrac{45}{x + 4} \times \dfrac{x}{x}= 6\) \(\dfrac{60(x + 4) + 45x}{(x + 4)x} = 6\) \(60x + 240 + 45x = 6x^2 + 24x\) \(6x^2 - 81x - 240 = 0\) \(2x^2 - 27x - 80 = 0\)